Biosphere 2 Research Wiki » » Blog categories » Data-logging and Controls

Data-logging and Controls

Last modified by Michael Sibayan on 2015/02/13 12:32
Comments (0) · Attachments (0) · History

Category: Data-logging and Controls (30 posts) [RSS]

Nov 14 2016

Seepage clogged

Seepage clogged and fixed,flow rates down while fixing ...

Nov 07 2016

Irrigation valve actuator swap 2016117093157

On Nov. 6, 2016 from about 10:45 PM - 11:30 PM Aaron and Michael performed work on the electric actuators controlling/monitoring LEO irrigation. 

Several of the actuators were not responding (No diagnostics have been performed), so we swapped around actuators such that each bay now should have a functional actuator on both the MULTI1 and MULTI2 circuit. 

Due to this change several irrigation circuits will display as "in transition" because they have been physically removed / electrically disconnected.

Aug 16 2016

SDI-12 sensors removed

Today several more SDI-12 sensors were removed from the bus due to large power draw in:
East bay zones P & D
Center bay zone K

LEO-C_4_0_2_HMP60 malfunction

The HMP60 in the center bay, mast 4,0 position 2 appears to be incorrectly reporting RH. Output signal has been verified and matches recorded value.

Jul 25 2016

LEO SDI-12 system powered on for cooling experiment

The SDI-12 systems in all three bays were powered on to support an air flow / cooling experiment. VWC is required for the surface heat flux computation.

Jul 24 2016

LEO Sampling Interval Change

July 24, 2016
The sampling interval for most LEO sensors, excluding 5TM, MPS-2, and CS451, has been set to 1 minute for a series of above slope air flow experiments. Since there is no anticipated irrigation I have decided to leave the SDI-12 sensors (listed above) at the standard 15 minute logging interval.

Jun 21 2016

Mini LEO data logger software update 2016621030609

At approximately 15:00 on Wednesday, June 21 the LEO Mini data logger software was updated and the logger was restarted.
This update is intended to address an issue with loss of precision of the large counter & volume values of the tipping bucket. The 5-digit level of precision has been removed for all double precision data types within the logging software.

Jun 09 2016

Mini LEO calibrations 20160609

Antonio has performed calibrations for the drainage tipping bucket and EC meter. As of today (06/09/2016) the scales for these two devices have been updated on the data logger.

TB - 0.007007 liters per tip
EC = 711.167 * voltage - 124.152

Aug 27 2015

LEO-Mini software update 20150827

Maintenance started at about 14:30 on August 27, 2015. The software on the Mini LEO data-logger (NI Compact RIO) is getting upgraded to the the 2014 model and the OS is getting upgraded to 14.0.1. Additional information will be posted as replies to this entry.

Jul 28 2015

LEO-E_4_0_1_HMP60 replaced

July 28, 2015 - Nate replaced the damaged HMP60 by cutting the old cable connector and splicing a new one. Data appears to be valid.

Created by Michael Sibayan on 2015/02/13 12:32

This wiki is licensed under a Creative Commons 2.0 license
XWiki Enterprise 6.0.1 - Documentation